In isobaric process a gas did A=270J of work. By how much will the volume increase in this process, if gas pressure p=4.5×10⁴pa
1
Expert's answer
2020-08-05T16:42:05-0400

In isobaric process the work a gas did can be expressed in the following way:


"A = p\Delta V"

where "\Delta V" is the amount by which the volume will increase in this process and "p = 4.5\times 10^4 Pa" is gas pressure. Expressing "\Delta V", obtain:


"\Delta V = \dfrac{A}{p} = \dfrac{270}{4.5\times 10^4} = 6\times 10^{-3} m^3"

Answer. "6\times 10^{-3} m^3".